# Energy Rank Alignment: Using Preference Optimization to Search Chemical Space at Scale

### v1

@@ 3 Theoretical Analysis
By construction, $\pi(\boldsymbol{y}|\boldsymbol{x})$ does not depend on $\boldsymbol{y}^{\prime}$ so the prefactor must be purely a function of $\boldsymbol{x}$, which completes the proof, using Lemma 3.1.
Gradients of $\mathcal{L}^{\textrm{ERA}}$.
[…]
4 Experiments
We test ERA on both chemical and language tasks to shed light on the following questions: 1) Can we use ERA to robustly fine-tune our model to generate samples according to a desired distribution? 2) What is the effect of changing the inverse-temperature $\beta$ during ERA? 3) Do we maintain sample diversity (and validity) without regularizing to remain close to a reference policy, and what is the effect of increased regularization? 4) Can we simultaneously target multiple properties with high fidelity, and how can we trade off between desired properties? 5) Can we carry out ERA on higher capacity models with “weak” signals from smaller models?
4.1 Generating molecules with desired properties
We use a decoder-only representation for the molecular generator [4], where the generator has 2 layers, an embedding dimension of 512, a vocabulary of 324 tokens, and totals 3.5M parameters. Starting from a random initialization, we carry out pretraining on a dataset of 2.4M small molecules from the ChEMBL database [37] for 180 epochs. This version of the model is not conditioned on a prompt and generates a small molecule given just a start-of-sequence token. We use this pretrained model as our reference policy for all unprompted molecular alignment tasks (Sec. 4.1.1). In Sec. 4.1.2, we generate molecules conditioned on a prompt using a generator that was trained to carry out sampling with a prompt molecule.
Central to ERA is, of course, access to a computable energy function. As a proof-of-concept, here we consider 5 different properties for which the corresponding energy function is easily evaluable: Quantitative Estimate of Drug-Likeness (QED) [6], Wildman-Crippen LogP (LogP) [36], Ring Count, Molar Refractivity (MR) [36], and Tanimoto Similarity [26]. Briefly, LogP is a measure of the hydrophobicity of a molecule, MR is a measure of the polarizability of the molecule, and Tanimoto similarity is a measure of the similarity between two molecules (see Appendix D.2).
4.1.1 Unprompted molecular alignment
First, we independently target four different properties using ERA with an unprompted molecular generator (Fig. 2). Using the reference policy, we generate a dataset $\mathcal{D}=\{\boldsymbol{y}_{1}^{(i)},\boldsymbol{y}_{2}^{(i)},U(\boldsymbol{y}_{1}^{(i)}),U(\boldsymbol{y}_{2}^{(i)})\}_{i=1}^{N}$ and carry out energy rank alignment on $\pi_{\boldsymbol{\theta}},$ where $\pi_{\boldsymbol{\theta}}$ is initialized using the weights of $\pi_{\rm ref}.$ Here, $\boldsymbol{y}_{1},\boldsymbol{y}_{2}\sim\pi_{\rm ref}$ and $\boldsymbol{y}$ and $U(\boldsymbol{y})$ denote the generated molecule and its corresponding energy, respectively. For MR, Ring Count, and LogP, we define the energy $U$ to be a harmonic potential centered at a target value. For QED, we define the energy to be the negative logarithm of QED and vary $\beta$ to assess its impact on alignment (see Table 12). In Fig. 2, we see that we successfully shift the distribution to target means that are both greater and lower than the average value of MR, Ring Count, and LogP under the reference policy. Furthermore, in the alignment of QED, we observe the effect of changing $\beta$ on the learned policy; with increased $\beta,$ the learned policy concentrates around low-energy samples (i.e. near $\textrm{QED}=1$), and with lower $\beta,$ the learned policy samples a greater range of QED values, as expected. We note that for each of these four experiments, we did not regularize towards the reference policy (i.e. $\gamma=0$). Even so, we were able to maintain both sample diversity and maintain appreciable sample validity (see Fig. 7 and Table 3).
Many molecular design tasks require balancing multiple properties, and designing an objective for multi-property alignment is straightforward within the ERA framework. To demonstrate this, we generate molecules with both high QED and LogP using ERA with an energy function weighted by property-specific $\beta$: $U=\beta_{\rm QED}U_{\rm QED}+\beta_{\rm LogP}U_{\rm LogP}$ (see Table 14 for details on energy function). We carry out ERA with different pairs of $(\beta_{\rm QED},\ \beta_{\rm LogP})$ using the same procedure as above, and from Fig. 3, we see that we target multiple properties with varying fidelity by simply modulating the value of property-specific $\beta.$ Ultimately, increasing the $\beta$ for an individual property enables us to favor higher values of that property in multi-property alignment setting. In this case, we also do not regularize with the KL-divergence to the reference policy and again maintain sample diversity and validity (see Fig. 8 and Table 4)
4.1.2 Prompted molecular alignment
Refer to captionRefer to caption
Figure 4: Prompted multi-property molecular generator alignment. From left to right: Tanimoto similarities computed between the prompt and sampled molecules for both aligned and unaligned policies (QED and Tanimoto alignment), per-prompt difference in the average QED under aligned and unaligned policies (QED and Tanimoto alignment), Tanimoto similarities computed between the prompt and sampled molecules for both aligned and unaligned policies (LogP and Tanimoto alignment), and per-prompt difference in the average LogP under aligned and unaligned policies (LogP and Tanimoto alignment). With alignment, we target higher QED and LogP values, while still sampling molecules chemically similar—but not identical—to prompt molecule.
Inspired by the task of lead optimization in drug discovery efforts [16], we ask whether we can use ERA to train a molecular generator that can sample a molecule that is both similar to the prompt molecule and also exhibits some desired property.
First, we fine-tune the pretrained molecular generator to enable prompted molecular generation (see Appendix D.3.2) and use this fine-tuned model as our reference policy for all prompted molecular alignment tasks. This reference policy disproportionately samples molecules that are identical (i.e. a Tanimoto similarity of 1.0) to the prompt molecule (see Fig. 4), so we carry out multi-property alignment on this reference policy to generate molecules that are similar—but not identical—to the prompt molecule and also have a high drug-likeness as measured by QED. Using ERA, we optimize the reference policy with a generated dataset $\mathcal{D}=\{(\boldsymbol{y}_{1}^{(i)},\boldsymbol{x}^{(i)}),(\boldsymbol{y}_{2}^{(i)},\boldsymbol{x}^{(i)}),U(\boldsymbol{y}_{1}^{(i)},\boldsymbol{x}^{(i)}),U(\boldsymbol{y}_{2}^{(i)},\boldsymbol{x}^{(i)})\}_{i=1}^{N},$ where we sample four molecules for each prompt molecule from the reference policy and consider all possible preference pairs for a total of six preference pairs per prompt molecule (see Appendix D.2 for full details on energy used).
We observe that the per-prompt average QED under the optimized policy for a given prompt is higher than the corresponding average under the reference policy (Fig. 4). Furthermore, we see that we are able to sample a diverse set of molecules that are chemically similar to the prompt molecule, and also chemically valid (see Figure 9, Table 5). We repeat the experiment with a related objective of generating molecules similar to the prompt molecule with a high LogP instead and again observe that we increase the per-prompt average LogP under the optimized policy relative to the reference policy without degrading sample diversity and validity. For both of these experiments, we required regularization to the reference policy. With no regularization, the aligned generator would almost exclusively sample sequences that were chemically invalid ($<25\%$ chemical validity). Finally, we note that the increases in QED and LogP in Fig. 4 are smaller relative to the increases in Fig. 2 because the samples are now conditioned to remain proximal to the prompt molecule, which restricts the chemical space that can be explored.
4.2 AI-guided alignment of large language models
Figure 5: AI-guided alignment of LLMs. Average sentiment of responses from aligned GPT-2 model across all prompts. (left). Proportion of unsafe content relative to unaligned model of responses aligned LLaMA2-13B model across all prompts (right). 5.4% of all responses from unaligned model were classified as unsafe. Error bars too small to be shown.
We test the generality of ERA by applying it to align large language models (LLMs). Similar to the experiments in [25], we first carry out ERA on a GPT-2 model [24] fine-tuned on movies reviews from IMDb [18]. We use a pretrained sentiment classifier [14] to evaluate the energies—where lower energies correspond to more positive sentiments—of sampled responses from the reference policy and carry out ERA using the same approach as in Section 4.1.2 (see Appendix E.1). We vary the regularization strength $\gamma$ and inverse-temperature $\beta$ on the average sentiment and observe that across all regularization strengths, with increasing $\beta,$ the average sentiment becomes more positive. Increasing regularization also elicits more positive sentiments. Qualitatively, with lower regularization, we observe that text quality degrades and becomes less coherent, likely resulting in lower average sentiment predictions by the sentiment model. Regularization here is important to ensure high quality text samples.
We next leverage a “weak” AI supervisor to carry out LLM alignment, a task sometimes called “superalignment” [8]. In the present context, we order “weak” vs. “strong” models based on their parameter count (within the same family) and empirical performance; i.e., LLaMA2-7B is weaker than LLaMA2-13B. Here, the weak model does not necessarily contain the complexity of the stronger model but can weakly discern between different outputs of a stronger model. Given a sample $\boldsymbol{y}_{i}\sim\pi_{\rm strong}(\boldsymbol{y}|\boldsymbol{x}),$ we define the energy using the weak model $U(\boldsymbol{y}_{i}|\boldsymbol{x})=-\log\pi_{\rm weak}(\boldsymbol{y}_{i}|\boldsymbol{x}).$
We test weak-to-strong alignment using a previously aligned LLaMA2-7B-Chat (meta-llama/Llama-2-7b-chat) to optimize an unaligned LLaMA2-13B (meta-llama/Llama-2-13b) model [33]. Using prompts from the Anthropic Helpful and Harmless dialogue dataset [5], we first carry out a short supervised fine-tuning step of LLaMA2-13B to ensure it can output text in a chat-like format (see Appendix E.2). Using this reference policy, we generate a dataset with energies computed from the smaller LLaMA2-7B-Chat model and carry out ERA as above, again across varying $\gamma$ and $\beta.$ We evaluate the “safety” of generated samples using Meta LLama Guard 2 (meta-llama/Meta-Llama-Guard-2-8B) [15]. We observe that as we increase $\beta,$ the proportion of unsafe content relative to the unaligned, reference model decreases, with over a 90% drop between the unaligned model and the models aligned with the highest $\beta$ across all $\gamma.$ For these experiments, we observe that varying regularization strengths has a minimal effect and that we are in fact able to generate coherent sentences with no regularization, with strong regularization hurting performance for $\beta=0.1.$ Finally, we compare ERA and DPO in Appendix E.2 and observe that with our implementation of DPO, we are able to generate lower energy samples, but that it is prone to mode collapse. We caution that our implementation of DPO is likely not optimal and that we did not exhaustively tune the hyperparameters of DPO due to resource constraints.
5 Conclusions and Limitations
This paper introduces energy rank alignment, a simple and effective algorithm for policy optimization with an explicit reward model. We find that ERA is stable without extensive hyperparameter tuning, and sufficiently general to successfully align both application-specific transformers for chemical search problems as well as generative pre-trained transformers for language. The algorithm exhibits strong performance with a variety of reward models, even ones with relatively weak signal, such as the AI feedback of LLaMA2-7B-Chat. Interestingly, with this approach we are able to reduce unsafe content by more than 90% with no human preference data.
We analyze the minimizers of the ERA objective and find that they differ from the minimizers of popular policy alignment algorithms DPO and PPO in an important way: unlike PPO, the strength of regularization to the reference policy that we add is controlled by a parameter $\gamma$, while the entropy of the target distribution is independently tuned by a distinct parameter $\beta$. This means that we can avoid greedy policies by keeping $\beta$ small—amplifying fluctuations around the optimum of the reward model $-U$—while reducing the influence of the reference policy by taking $\gamma$ small. Our objective leads to easily interpretable sample-wise gradients which highlight the importance of a reward model relative to DPO in the sampled objective. Similar observations about the inadequacy of the DPO objective for finite preference observations were also made theoretically in Azar et al. [3].
Limitations:
First, our approach requires a reward model, which can be difficult to train or design, especially for complex tasks. While we observed that ERA makes an appreciable impact even with weak supervision from an AI chat model, this sort of proxy may not be available for more complex tasks. For example, optimizing small molecules for high binding affinity to a target protein would require expensive and noisy evaluations of a reward model, which likely limits the scope of molecular design to problems where the reward can be computed somewhat efficiently. A second limitation of our present work is that we do not train the molecular transformer to favor synthetic accessibility nor do we explicitly seek to obtain molecules that are easily synthesized experimentally. There are models that seek to evaluate synthesizability computationally that could be used in our rewards, which we plan to explore in future work [11]. A final limitation of our current work is the moderate scale of our numerical experiments due to our limited compute resources, including the inadequate hyperparameter tuning for the DPO baseline for Fig. 5.
Appendix A Code Availability
Appendix B Detailed Theoretical Analysis
@@
Using Lemma  B.1 it is straightforward to prove the result in the main text Proposition 3.2. For completeness, we re-state that result here and refer the reader to the main text for the complete argument.
Proposition B.2
@@
B.1 Comparison with PPO Objective
@@
B.2 Comparison with DPO Objective
@@
Appendix C ERA implementation
@@
Appendix D Details for molecular generator experiments
D.1 Pretraining details
@@
Our trained molecular generator consisted of just the encoder block of a standard multi-head attention transformer [34]. Finally, the model had 2 layers, 8 heads, and a width of 512. For pretraining, we used an Adam optimizer with a learning rate of $1.0*10^{-5}$. We emphasize that this pretrained generator samples molecules in an unprompted fashion; given just a start-of-sequence token, we can autoregressively generate a sequence of tokens. Moreover, it is possible that this sequence of tokens corresponds to a molecule that is not chemically valid, and we find that around 88% of all generated molecules are chemically valid. Lastly, we measure the diversity of the pretrained molecular generator by first generating 1500 molecules and then computing the Tanimoto similarity between every pair of molecules. We plot the distribution of all pairwise Tanimoto similarities from this sample and from all pariwise Tanimoto similarities from 1500 randomly sampled molecules from the original dataset in Fig. 6. We observe that we can generate molecules that are quite distinct (i.e. low Tanimoto similarity) in comparison with all other molecules.
Figure 6: Chemical diversity of samples from training dataset and from unprompted molecular generator (unaligned) as measured by pairwise Tanimoto similarities. Lower Tanimoto similarities correspond to more chemically dissimilar molecules.
D.2 Chemical properties
We investigated aligning the molecule generator to several target chemical properties, which we detail below. All of the properties can be easily computed using the RDKit package. We list the energy function and parameters used for the corresponding energy functions for each of these properties in Table 1.
@@ D.2 Chemical properties
Table 1: Definitions of energy functions (in reduced units) used for each of the five chemical properties investigated in this work. Here $\boldsymbol{y}$ refers to the generated molecule.
Under the definitions of the energy functions in Table 1, it is possible for a generated sequence to not be chemically valid. For these cases, we manually define energies that are sufficiently high to penalize that outcome and we report these values in Table 2. Furthermore, when the computed QED or Tanimoto Similarity is 0, the energy is infinite, and to ensure numerical stability, we set the value of the energies to be 4.5 and 10 respectively. Finally, in the prompted molecular generator experiments in Section 4.1.2, we assign an energy of 3.5 to the setting where Tanimoto similarity between the generated and prompt molecule is 1.0 (i.e they are the same) in order to penalize this outcome. Here, all energy and $\beta$ values are reported in reduced units.
@@
Table 2: Property-specific energy values (in reduced units) used to treat chemically invalid sequences.
D.3 Molecular alignment details
D.3.1 Unprompted molecular generation
Figure 7: Chemical diversity of samples from unprompted molecular generator after alignment as measured by pairwise Tanimoto similarities. (See Fig. 2, Section 4.1.1)
@@ D.3.1 Unprompted molecular generation
Table 3: Percentage of generated sequences that were chemically valid for samples from unprompted molecular generator after alignment. (See Fig. 2, Section 4.1.1).
Figure 8: Chemical diversity of samples from unprompted molecular generator after multi-property alignment as measured by pairwise Tanimoto similarities. (See Fig. 3, Section 4.1.1).
@@
Table 4: Percentage of generated sequences that were chemically valid for samples from unprompted molecular generator after multi-property alignment. (See Fig. 3, Section 4.1.1).
We first investigated aligning the unprompted molecular generator to sample small-molecules with desired properties. We carried out alignment using the property-specific energies described in Table 1. All alignment properties were initialized with the weights of the pretrained model and trained using an Adam optimizer with learning rate $1.0*10^{-6}$. We tabulate the chemical validity for single-property alignment in Table 3 and for multi-property alignment in Table 4. While we do see a drop in chemical validity after alignment, we see that a majority of the samples we generate post-alignment are still chemically valid despite no regularization to a reference policy. We measure the chemical diversity for these experiments by computing all pairwise Tanimoto similarities from all chemically valid predictions of 1500 generated molecules. We visualize the chemical diversity for single-property experiments in Fig. 7 and multi-property experiments in Fig. 8. We observe that the samples are still highly diverse chemically after alignment. All plots in Fig. 2 and Fig. 3 were computed using 1500 generated molecules per experiment.
D.3.2 Prompted molecular generation
Refer to captionRefer to caption
Figure 9: Sample molecules from prompted molecular generator after multi-property alignment experiments: QED and Tanimoto (left) and LogP and Tanimoto (right). With alignment, generated molecules are diverse, while still chemically similar to prompt molecule.

### v2

@@ Appendix B Theoretical Analysis
By construction, $\pi(\boldsymbol{y}|\boldsymbol{x})$ does not depend on $\boldsymbol{y}^{\prime}$ so the prefactor must be purely a function of $\boldsymbol{x}$, which completes the proof, using Lemma B.1.
Appendix C Detailed Theoretical Analysis
@@
Using Lemma  C.1 it is straightforward to prove the result in Proposition B.2. For completeness, we re-state that result here and refer the reader to Appendix B for the complete argument.
Proposition C.2
@@
C.1 Loss functions for $\pi_{\boldsymbol{\theta}}$:
Proximal Policy Optimization (PPO) optimizes an indirect, proximal objective to minimize an objective closely related to (2) (cf. Appendix C). Direct Preference Optimization (DPO) treats the negative reward function $U$ implicitly and directly maximizes the likelihood of $p(\boldsymbol{y}\succ\boldsymbol{y}^{\prime}|\boldsymbol{x})$. Our objectives differ from both approaches: like DPO, we directly optimize the policy using an explicit, gradient-based objective, but, in contrast, we use a reward function directly in our objective. The losses we build are thus amenable to both offline (samples from $\pi_{\rm ref}$) and online (samples from $\pi_{\boldsymbol{\theta}}$) policy alignment, as explained below. Choosing to optimize the objective online has been shown to have important consequences on performance [38], though we focus here on the setting where samples are drawn offline.
where
instead of (9). One issue that arises with this scheme is differentiation with respect to the parameters of the policy $\boldsymbol{\theta}$ because $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ are decoded into discrete tokens, an operation that is not differentiable. To remedy this, we importance sample with a reference policy
C.2 Gradients of $\mathcal{L}^{\textrm{ERA}}$.
[…]
C.3 Comparison with PPO Objective
@@
C.4 Comparison with DPO Objective
@@
Figure S6: Comparison of DPO ($\beta_{\rm DPO}$=0.1) and ERA ($\beta=20.0$ and $\gamma=0.0$) on the task of maximizing QED (Section 3.1.1) a) Chemical validity degrades significantly for DPO and slightly degrades for ERA as a function of checkpoints. b) At each checkpoint, 20k molecules are sampled and mean QED of valid molecules plotted with shading corresponding to 1 standard deviation. Both DPO and ERA alignment runs were trained with the same hyperparameters and dataset, with checkpoints saved every 100 epochs.
We carry out an online evaluation of DPO and ERA on the task of generating small-molecules with high QED, as in Figure 2. We align DPO using the same dataset and hyperparameters as ERA with $\beta_{DPO}=0.1$ and train for thousands of checkpoints (over 72 GPU-hours). We load intermediate checkpoints for both the DPO and ERA ($\beta_{ERA}=20.0$, $\gamma_{ERA}=0.0$) runs and carry out inference (Figure S6). We observe that at the first saved checkpoint of the DPO alignment run, the model generates molecules with high QED scores but with low validity ($\sim$20%). However, upon further training, the chemical validity of further checkpoints drops to 0% for the remaining runs, despite the overall DPO training and validation losses still dropping.
With ERA, we see that we are able to similarly sample high QED small-molecules with reasonably high chemical validity ($\sim$85%). While the validity does drop over subsequent checkpoints, it does not do so precipitously. Moreover, the ERA-based alignment had no regularization ($\gamma=0$), and in the main text, we document how having non-zero $\gamma$ can enable increases in chemical validity (Figure S10). Finally, we note that we did not extensively tune the hyperparameters for DPO, and it is possible that a different set of hyperparameters would elicit a more desired outcome; however, the lack of meaningful regularization in DPO [3] and its performance degradation in online metrics has been well-documented [10].
Appendix D ERA implementation
@@
Appendix E Details for molecular generator experiments
E.1 Pretraining details
@@
Our trained molecular generator consisted of just the encoder block of a standard multi-head attention transformer [39]. Finally, the model had 2 layers, 8 heads, and a width of 512. For pretraining, we used an Adam optimizer with a learning rate of $1.0*10^{-5}$. We emphasize that this pretrained generator samples molecules in an unprompted fashion; given just a start-of-sequence token, we can autoregressively generate a sequence of tokens. Moreover, it is possible that this sequence of tokens corresponds to a molecule that is not chemically valid, and we find that around 88% of all generated molecules are chemically valid. Lastly, we measure the diversity of the pretrained molecular generator by first generating 1500 molecules and then computing the Tanimoto similarity between every pair of molecules. We plot the distribution of all pairwise Tanimoto similarities from this sample and from all pariwise Tanimoto similarities from 1500 randomly sampled molecules from the original dataset in Fig. S7. We observe that we can generate molecules that are quite distinct (i.e. low Tanimoto similarity) in comparison with all other molecules.
Figure S7: Chemical diversity of samples from training dataset and from unprompted molecular generator (unaligned) as measured by pairwise Tanimoto similarities. Lower Tanimoto similarities correspond to more chemically dissimilar molecules.
E.2 Chemical properties
We investigated aligning the molecule generator to several target chemical properties, which we detail below. All of the properties can be easily computed using either the RDKit package or the tdc [20] package. We list the energy function and parameters used for the corresponding energy functions for each of these properties in Table S2.
@@ E.2 Chemical properties
Table S2: Definitions of energy functions (in reduced units) used for each of the five chemical properties investigated in this work. Here $\boldsymbol{y}$ refers to the generated molecule.
Under the definitions of the energy functions in Table S2, it is possible for a generated sequence to not be chemically valid. For these cases, we manually define energies that are sufficiently high to penalize that outcome and we report these values in Table S3. Furthermore, when the computed QED or Tanimoto Similarity is 0, the energy is infinite, and to ensure numerical stability, we set the value of the energies to be 4.5 and 10 respectively. Finally, in the prompted molecular generator experiments in Section 3.1.2, we assign an energy of 3.5 to the setting where Tanimoto similarity between the generated and prompt molecule is 1.0 (i.e they are the same) in order to penalize this outcome. Here, all energy and $\beta$ values are reported in reduced units.
@@
Table S3: Property-specific energy values (in reduced units) used to treat chemically invalid sequences.
E.3 Molecular alignment details
E.3.1 Unprompted molecular generation (RDKit oracles)
Figure S8: Chemical diversity of samples from unprompted molecular generator after alignment as measured by pairwise Tanimoto similarities. (See Fig. 2, Section 3.1.1)
@@ E.3.1 Unprompted molecular generation (RDKit oracles)
Table S4: Percentage of generated sequences that were chemically valid for samples from unprompted molecular generator after alignment. (See Fig. 2, Section 3.1.1).
Figure S9: Chemical diversity of samples from unprompted molecular generator after multi-property alignment as measured by pairwise Tanimoto similarities. (See Fig. 3, Section 3.1.1).
@@
Table S5: Percentage of generated sequences that were chemically valid for samples from unprompted molecular generator after multi-property alignment. (See Fig. 3, Section 3.1.1).
We first investigated aligning the unprompted molecular generator to sample small-molecules with desired properties. We carried out alignment using the property-specific energies described in Table S2. All alignment properties were initialized with the weights of the pretrained model and trained using an Adam optimizer with learning rate $1.0*10^{-6}$. We tabulate the chemical validity for single-property alignment in Table S4 and for multi-property alignment in Table S5. While we do see a drop in chemical validity after alignment, we see that a majority of the samples we generate post-alignment are still chemically valid despite no regularization to a reference policy. We measure the chemical diversity for these experiments by computing all pairwise Tanimoto similarities from all chemically valid predictions of 1500 generated molecules. We visualize the chemical diversity for single-property experiments in Fig. S8 and multi-property experiments in Fig. S9. We observe that the samples are still highly diverse chemically after alignment. All plots in Fig. 2 and Fig. 3 were computed using 1500 generated molecules per experiment.
E.3.2 Prompted molecular generation (RDKit oracles)
Figure S10: Prompted multi-property molecular generator alignment. From left to right: Tanimoto similarities computed between the prompt and sampled molecules for both aligned and unaligned policies (QED and Tanimoto alignment), per-prompt difference in the average QED under aligned and unaligned policies (QED and Tanimoto alignment), Tanimoto similarities computed between the prompt and sampled molecules for both aligned and unaligned policies (LogP and Tanimoto alignment), and per-prompt difference in the average LogP under aligned and unaligned policies (LogP and Tanimoto alignment). With alignment, we target higher QED and LogP values, while still sampling molecules chemically similar—but not identical to—the prompt molecule.
Figure S11: Sample molecules from prompted molecular generator after multi-property alignment experiments: QED and Tanimoto (left) and LogP and Tanimoto (right). With alignment, generated molecules are diverse, while still chemically similar to prompt molecule.